Comparative data from the statement of financial position of Benton Ltd. are shown below.
2018 2017 2016
Current assets $1,524,000 $1,182,000 $1,218,000
Property, plant, and equipment 3,155,000 2,786,000 2,843,000
Goodwill 83,000 93,000 -0-
    Total assets $4,762,000 $4,061,000 $4,061,000
Using horizontal analysis, calculate the percentage of the base-year amount, using 2016 as the base year. (Round answers to 1 decimal place, e.g. 52.7%.)
2018 2017 2016
Current assets

%

%

%
Property, plant, and equipment

%

%

%
     Total assets

%

%

%
Using vertical analysis, calculate the percentage of the base amount for each year. (Round percentages to 1 decimal places, e.g. 52.7%.)
2018 2017 2016
Amount Percentage Amount Percentage Amount Percentage
Current assets $

% $

% $

%
Property, plant, and equipment

%

%

%
Goodwill

%

%

%
Total assets $

% $

% $

%
Answer #1

Answers

[1]

2018

2017

2016

Current Assets

125.1%

97.0%

100.0%

PPE

111.0%

98.0%

100.0%

Goodwill

0.0%

0.0%

0.0%

Total Assets

117.3%

100.0%

100.0%

--Working

2018

2017

2016

Current Assets

=1524000/1218000

=1182000/1218000

=1218000/1218000

PPE

=3155000/2843000

=2786000/2843000

=2843000/2843000

Goodwill

0

0

0

Total Assets

=4762000/4061000

=4061000/4061000

=4061000/4061000

[2]

2018

2017

2016

Current Assets

$1,524,000

32.0%

$1,182,000

29.1%

$1,218,000

30.0%

PPE

$3,155,000

66.3%

$2,786,000

68.6%

$2,843,000

70.0%

Goodwill

$83,000

1.7%

$93,000

2.3%

$0

0.0%

Total Assets

$4,762,000

100.0%

$4,061,000

100.0%

$4,061,000

100.0%